### v1

@@ Abstract
Large Vision-Language Models (LVLMs) have shown significant progress in well responding to visual-instructions from users. However, these instructions, encompassing images and text, are susceptible to both intentional and inadvertent attacks. Despite the critical importance of LVLMs’ robustness against such threats, current research in this area remains limited. To bridge this gap, we introduce AVIBench, a framework designed to analyze the robustness of LVLMs when facing various adversarial visual-instructions (AVIs), including four types of image-based AVIs, ten types of text-based AVIs, and nine types of content bias AVIs (such as gender, violence, cultural, and racial biases, among others). We generate 260K AVIs encompassing five categories of multimodal capabilities (nine tasks) and content bias. We then conduct a comprehensive evaluation involving 14 open-source LVLMs to assess their performance. AVIBench also serves as a convenient tool for practitioners to evaluate the robustness of LVLMs against AVIs. Our findings and extensive experimental results shed light on the vulnerabilities of LVLMs, and highlight that inherent biases exist even in advanced closed-source LVLMs like GeminiProVision and GPT-4V. This underscores the importance of enhancing the robustness, security, and fairness of LVLMs. The source code and benchmark will be made publicly available. * indicates corresponding author.
Keywords:
@@
Large Vision-Language Model Adversarial Visual-Instructions Bias Evaluation
1 Introduction
@@
Over the past year, Large Language Models (LLMs) have achieved significant milestones, consistently demonstrating exceptional performance across a diverse range of natural language processing tasks. This success has spurred the development of LLM-based applications, reshaping our daily lives. More recently, alongside advanced closed-source Large Vision-Language Models (LVLMs) like GeminiProVision [1] and GPT-4V(ision) [2], many open-source LVLMs have emerged, such as Otter [3], InternLM-XComposer [4], ShareGPT4V [5], and Moe-LLaVA [6]. These open-source models propose various architectures and training methods to enhance the capabilities of powerful LLMs like Vicuna [7] and LLaMA [8], enabling them to understand images and perform multimodal tasks such as visual question answering [9], multimodal conversation [10], and complex scene comprehension [3]. Considering that LVLMs form the foundation for next-generation AI applications, addressing concerns related to their robustness, security, and bias is of utmost importance.
Figure 1: The overview of AVIBench.
@@
LVLMs employ two input modalities, text and images, both susceptible to adversarial perturbations. While pioneering studies [11, 12] have assessed LLMs’ robustness against text-based attacks, there is a lack of specific exploration targeting LVLMs. Recent investigations of image attacks have examined limited LVLMs’ resilience against white-box attacks [13, 14], backdoor attacks [15], query-based black-box attacks [16], and transfer-based black-box attacks [17]. However, transfer-based black-box attacks rely on surrogate models to execute the attacks, posing challenges in finding an LVLM-agnostic surrogate model applicable to all LVLMs. White-box attacks, backdoor attacks, and query-based black-box attacks, which depend on the output probability distributions of LVLMs, may be impractical for online-accessed models, particularly closed-source LVLMs. Moreover, these attack methods may be constrained by their specific task design, such as image captioning [18] or visual question answering [13], thus limiting the evaluation’s comprehensiveness.
Figure 2: Comparison of LVLMs’ robustness of adversarial visual-instructions. The robustness of adversarial visual-instructions about (a) image corruptions, (b) decision-based optimized black-box image attacks, (c) text attacks, and (d) content bias attacks.
@@
This paper introduces AVIBench, a comprehensive benchmark designed to evaluate LVLMs’ robustness in the face of adversarial visual-instructions (text-image pairs), as illustrated in Fig. 1. AVIBench encompasses a diverse set of adversarial visual-instructions (AVIs) that target text and images. Specifically, we adapt LVLM-agnostic and output probability distributions-agnostic black-box attack methods to target LVLMs, resulting in a total of 10 types of text-based AVIs and 4 types of image-based AVIs. In addition to these attacks, we also introduce 9 types of content bias AVIs, addressing issues related to gender, violence, culture, racial biases, and more, to evaluate the biases inherent in LVLMs comprehensively. AVIBench is mainly constructed from Tiny LVLM-eHub [19], which is a benchmark for five categories of multimodal capabilities (nine tasks). Finally, we construct 260K AVIs for AVIBench, which, along with our open-source code, can be utilized as a convenient tool to evaluate LVLMs’ defense against AVIs. We evaluate a total of 14 different open-source LVLMs using AVIBench and present the results in Fig. 2. Additionally, we evaluate closed-source LVLMs, including advanced systems like GeminiProVision and GPT-4V, using content bias AVIs.
Through extensive experimentation and analysis of the evaluation results, we make several noteworthy findings (detailed in Section 4). This paper serves a dual purpose by establishing a significant benchmark for assessing the robustness of LVLMs and potentially inspiring the development of mitigation and defense methodologies within the research community. Our main contributions can be summarized as follows:
  • We introduce AVIBench, a pioneering framework and versatile tool for evaluating the robustness of LVLMs on AVIs. AVIBench is designed to accommodate various tasks, models, and scenarios.
  • AVIBench generates a comprehensive dataset of 260K AVIs spanning five multimodal capabilities and content biases. This extensive dataset serves as a stringent evaluation benchmark, systematically probing LVLMs’ defense mechanisms against AVIs.
  • We evaluate the abilities of 14 open-source LVLMs to resist adversarial AVIs and show extensive experimental results and findings, which also offer convenience for developing robust LVLMs.
@@
2 Related Work
@@
Large Vision-Language Models. LVLMs like Otter [3], InstructBLIP [20], PandaGPT [21], InternLM-XComposer [4], LLaVA-1.5 [22], ShareGPT4V [5], and Moe-LLaVA [6] have made significant progress in multimodal tasks. These models align visual features with textual information by leveraging knowledge from LLMs like Vicuna [7] and LLaMA [8]. Techniques such as cross-attention layers [23], Q-Former [24], one project layer [25], and LoRA [26] have been used to bridge the gap between language and vision. Given the crucial role of LVLMs in future user interfaces and multimedia systems, ensuring their robustness, security, and fairness is paramount. Our paper focuses on evaluating LVLMs’ resilience against AVIs.
@@
Evaluation of Large Vision-Language Models. Recent advancements in LVLMs have led to improvements in datasets and evaluation methods. LVLM-eHub [27] and Tiny LVLM-eHub [19] organize multiple vision-language benchmarks, while MME Bench [28] introduces a new evaluation dataset. Other benchmarks like LAMM [29], MMBench [30], Seed Bench [31] also contribute to LVLM evaluation. However, there is a lack of research on assessing LVLMs’ resistance to attacks on text and image modalities. To address this, we propose a benchmark that evaluates LVLMs’ ability to withstand attacks on images, text, and bias.
@@
Attacks for Large Vision-Language Models. Pioneering research explored various LVLM-specific image attacks on limited LVLMs, including white-box attacks [13, 14], backdoor attacks [15], query-based black-box attacks [16], and transfer-based black-box attacks [17]. However, white-box attacks, backdoor attacks, and query-based black-box attacks require knowledge of the model’s output probability distribution; transfer-based black-box attacks require finding a surrogate model that is difficult to obtain for all LVLMs. Thus, we are the first to adapt LVLM-agnostic and output probability distribution-agnostic decision-based optimized image attacks specifically tailored for LVLMs. We also incorporate image corruption as an attack method, covering 5 multimodal capabilities across 9 subtasks, while Zhang et al. [32] only utilize image corruptions to evaluate limited LVLMs on the image captioning task. To target the text inputs of LVLMs, we draw inspiration from black-box text attacks originally designed for LLMs[11], adapting and expanding them for LVLMs. Furthermore, although previous studies have explored gender bias in the outputs of LVLMs [33, 34], we reveal the inherent biases that exist in LVLMs by building more comprehensive content bias AVIs.
3 AVIBench
@@
In this section, we first introduce the definition of Adversarial Visual-Instructions (AVIs), followed by an introduction to the key components of AVIBench: models, dataset, and the construction of AVIs.
3.1 Definition of Adversarial Visual-Instructions
@@
Unlike adversarial examples that cause "misclassification," Adversarial Visual-Instructions (AVIs) contain intentionally designed images and text that specifically manipulate LVLMs’ behavior in a broader sense. AVIs are intentionally crafted by adversaries to induce incorrect, unsafe, and harmful behavior in LVLMs, aligning with the broader definition of adversarial examples in [35][11].
3.2 Models
@@
We collect a total of 16 LVLMs, including 14 open-source models and 2 closed-source models, to create a model hub for evaluation. The open-source models consist of BLIP2 [24], LLaVA [36], MiniGPT-4 [25], mPLUG-owl [37], LLaMA-Adapter V2 [9], VPGTrans [38], Otter [3], InstructBLIP [20], PandaGPT [21], OpenFlamingo-V2 [23], InternLM-XComposer [4], LLaVA-1.5 [22], ShareGPT4V [5], and Moe-LLaVA [6]. The closed-source models are GeminiProVision [1] and GPT-4V(ision) [2]. To ensure a fair comparison, we carefully select the open-source LVLM versions with closely aligned parameters level. Details of these models can be found in the supplementary material.
3.3 Dataset
Our base dataset is derived from Tiny LVLM-eHub [19], which consists of 2,550 instructions and corresponding answers, and is organized into nine tasks that evaluate LVLM’s five multimodal capabilities: (1) Visual perception involves interpreting and understanding visual information. (2) Visual knowledge acquisition refers to the capability of LVLMs to acquire and understand visual information from images. (3) Visual reasoning involves the ability to reason and answer questions. (4) Visual commonsense measures the model’s comprehension of shared human knowledge about visual concepts. (5) Object hallucination refers to the phenomenon where LVLMs generate content that does not match the actual objects present in a given image.

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) have shown significant progress in responding well to visual-instructions from users. However, these instructions, encompassing images and text, are susceptible to both intentional and inadvertent attacks. Despite the critical importance of LVLMs’ robustness against such threats, current research in this area remains limited. To bridge this gap, we introduce B-AVIBench, a framework designed to analyze the robustness of LVLMs when facing various Black-box Adversarial Visual-Instructions (B-AVIs), including four types of image-based B-AVIs, ten types of text-based B-AVIs, and nine types of content bias B-AVIs (such as gender, violence, cultural, and racial biases, among others). We generate 316K B-AVIs encompassing five categories of multimodal capabilities (ten tasks) and content bias. We then conduct a comprehensive evaluation involving 14 open-source LVLMs to assess their performance. B-AVIBench also serves as a convenient tool for practitioners to evaluate the robustness of LVLMs against B-AVIs. Our findings and extensive experimental results shed light on the vulnerabilities of LVLMs, and highlight that inherent biases exist even in advanced closed-source LVLMs like GeminiProVision and GPT-4V. This underscores the importance of enhancing the robustness, security, and fairness of LVLMs. The source code and benchmark are available at https://github.com/zhanghao5201/B-AVIBench.
Index Terms:
@@
Large Vision-Language Model, Black-box, Adversarial Visual-Instructions, Bias Evaluation.
I Introduction
@@
Over the past year, Large Language Models (LLMs) have achieved significant milestones, consistently demonstrating exceptional performance across a diverse range of natural language processing tasks. This success has spurred the development of LLM-based applications [1], reshaping our daily lives. More recently, alongside advanced closed-source Large Vision-Language Models (LVLMs) like GeminiProVision [2] and GPT-4V(ision) [3], many open-source LVLMs have emerged, such as Otter [4], InternLM-XComposer [5], ShareGPT4V [6], and Moe-LLaVA [7]. These open-source models propose various architectures and training methods to enhance the capabilities of powerful LLMs like Vicuna [8] and LLaMA [9], enabling them to understand images and perform multimodal tasks such as visual question answering [10], multimodal conversation [11], and complex scene comprehension [4]. Considering that LVLMs form the foundation for next-generation AI applications [12, 13, 14], addressing concerns related to their robustness, security, and bias is of utmost importance.
Figure 1: The overview of B-AVIBench.
@@
LVLMs employ two input modalities, text and image, both susceptible to adversarial perturbations [15, 16, 17]. While pioneering studies [18, 19, 20] have assessed LLMs’ robustness against text-based attacks, there is a lack of specific exploration targeting LVLMs. Recent investigations of image attacks have examined limited LVLMs’ resilience against white-box attacks [21, 22], backdoor attacks [23], query-based black-box attacks [24], and transfer-based black-box attacks [25]. However, transfer-based black-box attacks rely on surrogate models to execute the attacks, posing challenges in finding an LVLM-agnostic surrogate model applicable to all LVLMs. White-box attacks, backdoor attacks, and query-based black-box attacks, which depend on the output probability distributions of LVLMs, may be impractical for online-accessed models, particularly closed-source LVLMs. Moreover, these attack methods may be constrained by their specific task design, such as image captioning [26] or visual question answering [21], thus limiting the evaluation’s comprehensiveness.
Figure 2: Comparison of LVLMs’ robustness scores of black-box adversarial visual-instructions for each LVLM. In each subfigure, we list the five most robust LVLMs under the corresponding attack, with the number inside the triangle indicating the rank. The definition of the robustness score is shown in Section IV-A.
@@
This paper introduces B-AVIBench, a comprehensive benchmark designed to evaluate LVLMs’ robustness in the face of black-box adversarial visual-instructions (text-image pairs), as illustrated in Fig. 1. B-AVIBench encompasses a diverse set of black-box adversarial visual-instructions (B-AVIs) that target text and images. Specifically, we adapt LVLM-agnostic and output probability distributions-agnostic black-box attack methods to target LVLMs, resulting in a total of 10 types of text-based B-AVIs and 4 types of image-based B-AVIs. In addition to these attacks, we also introduce 9 types of content bias B-AVIs, addressing issues related to gender, violence, culture, racial biases, and more, to evaluate the biases inherent in LVLMs comprehensively. B-AVIBench is mainly constructed from Tiny LVLM-eHub [27], which is a benchmark for five categories of multimodal capabilities (ten tasks). Finally, we construct 316K B-AVIs for B-AVIBench, which, along with our open-source code, can be utilized as a convenient tool to evaluate LVLMs’ defense against B-AVIs. We evaluate a total of 14 different open-source LVLMs using B-AVIBench and present the results in Fig. 2. Additionally, we evaluate closed-source LVLMs, including advanced systems like GeminiProVision and GPT-4V, using content bias B-AVIs.
Through extensive experimentation and analysis of the evaluation results, we make several noteworthy findings (detailed in Section IV). This paper serves a dual purpose by establishing a significant benchmark for assessing the robustness of LVLMs and potentially inspiring the development of mitigation and defense methodologies within the research community. Our main contributions can be summarized as follows:
  • We introduce B-AVIBench, a pioneering framework and versatile tool for evaluating the robustness of LVLMs on B-AVIs. B-AVIBench is designed to accommodate various tasks, models, and scenarios.
  • B-AVIBench generates a comprehensive dataset of 316K AVIs spanning five multimodal capabilities and content biases. This extensive dataset serves as a stringent evaluation benchmark, systematically probing LVLMs’ defense mechanisms against B-AVIs.
  • We evaluate the abilities of 14 open-source LVLMs to resist adversarial B-AVIs and show extensive experimental results and findings, which also offer convenience for developing robust LVLMs.
@@
The definitions of the abbreviations are as follows: In-BL., LA-V2, MGPT, m-owl, PGPT, VPGT, OF-2, In-XC., L-1.5, SGPT, Moe represent InstructBLIP, LLaMA-Adapter V2, MiniGPT-4, mPLUG-owl, PandaGPT, VPGTrans, OpenFlamingo-V2, InternLM-XComposer, LLaVA-1.5, ShareGPT4V, Moe-LLaVA, respectively. ‘VE’, ‘Adapter’, ‘ToP’, ‘TuP’, and ‘FC’ represent the vision encoder, adaptation module, total parameters of LLM, tuning parameters, and fully-connected layer, respectively. CC*, CC, VG, CY, L400, LC, QA*, SBU, ChatGPT, and LLaVA-I are consistent with the definition in LVLM-eHub [27]. C-name, Gau., and Imp. refer to corruption name, Gaussian, and Impulse respectively. Per., Kno., Rea., Com., and Hal. represent Visual Perception, Visual Knowledge Acquisition, Visual Reasoning, Visual Commonsense, and Object Hallucination respectively. P, B, and S correspond to PAR, Boundary, and SurFree respectively. Cha., Wor., Sen., Sem., Dee.ug, Inp.on represent character-level, word-level, sentence-level, and semantic-level, DeepWordBug, Input-reduction, respectively.
II Related Work
II-A Large Vision-Language Models
@@
LVLMs like Otter [4], InstructBLIP [28], PandaGPT [29], InternLM-XComposer [5], LLaVA-1.5 [30], ShareGPT4V [6], and Moe-LLaVA [7] have made significant progress in multimodal tasks. These models align visual features with textual information by leveraging knowledge from LLMs like Vicuna [8] and LLaMA [9]. Techniques such as cross-attention layers [31], Q-Former [32], one project layer [33], and LoRA [34] have been used to bridge the gap between language and vision. Given the crucial role of LVLMs in future user interfaces and multimedia systems, ensuring their robustness, security, and fairness is paramount. Our paper focuses on evaluating LVLMs’ resilience against B-AVIs.
II-B Evaluation of Large Vision-Language Models
@@
Recent advancements in LVLMs have led to improvements in datasets and evaluation methods. LVLM-eHub [35] and Tiny LVLM-eHub [27] organize multiple vision-language benchmarks, while MME Bench [36] introduces a new evaluation dataset. Other benchmarks like LAMM [37], MMBench [38], Seed Bench [39] also contribute to LVLM evaluation. However, there is a lack of research on assessing LVLMs’ resistance to attacks on text and image modalities. To address this, we propose a benchmark that evaluates LVLMs’ ability to withstand attacks on images, text, and bias.
II-C Attacks for Large Vision-Language Models
@@
In earlier studies, some works [40, 41] focus on the adversarial robustness of pre-trained vision-language models (VLMs), such as CLIP [42], when adapting to downstream datasets in the context of white-box attacks. They explore how to design adapters to enhance the robustness of VLMs for downstream tasks as a form of defense. However, the mechanisms and structures of the VLMs they studied differ significantly from those of LVLMs. VLMs utilize a contrastive learning mechanism, where the vision encoder and text encoder operate independently. In contrast, LVLMs employ a joint encoding mechanism based on the next-token prediction mechanism. Although some efforts have been made to help LVLMs resist attacks, such as FARE [43], which introduces unsupervised adversarial fine-tuning to defend against white-box attacks on CLIP-based LVLMs, and the approach in [22], which suggests using the Moderation API to filter out harmful instructions and outputs, these defense methods are specifically tailored to their respective attack strategies. As a result, diverse attack methods continue to pose significant threats to LVLMs. Specifically, pioneering research explored various LVLM-specific image attacks on limited LVLMs, including white-box attacks [21, 22, 44, 45, 46], backdoor attacks [23], query-based black-box attacks [24], and transfer-based black-box attacks [25]. However, white-box attacks, backdoor attacks, and query-based black-box attacks require knowledge of the model’s output probability distribution; transfer-based black-box attacks require finding a surrogate model that is difficult to obtain for all LVLMs. Thus, we are the first to adapt LVLM-agnostic and output probability distribution-agnostic decision-based optimized image attacks specifically tailored for LVLMs. We also incorporate image corruption as an attack method, covering 5 multimodal capabilities across 10 subtasks, while Zhang et al. [47] only utilize image corruptions to evaluate limited LVLMs on the image captioning task. To target the text inputs of LVLMs, we draw inspiration from black-box text attacks originally designed for LLMs[18], adapting and expanding them for LVLMs. Furthermore, although previous studies have explored gender bias in the outputs of LVLMs [48, 49], we reveal the inherent biases that exist in LVLMs by building more comprehensive content bias B-AVIs.
III B-AVIBench
@@
In this section, we first introduce the definition of Black-box Adversarial Visual-Instructions (B-AVIs), followed by an introduction to the key components of B-AVIBench: models, dataset, and the construction of B-AVIs.
III-A Definition of Black-box Adversarial Visual-Instructions
@@
Unlike adversarial examples that cause “misclassification,” Adversarial Visual-Instructions (AVIs) contain intentionally designed images and texts that specifically manipulate LVLMs’ behavior in a broader sense. AVIs are intentionally crafted by adversaries to induce incorrect, unsafe, and harmful behavior in LVLMs, aligning with the broader definition of adversarial examples in [50][18]. Black-box Adversarial Visual Instructions (B-AVIs) represent that the AVIs are built using black-box attack techniques.
TABLE I: Model configurations and data configurations of the LVLMs. The symbol indicates that the model is frozen. The composition of other data is described in their respective papers.
Model Model Configuration Image-Text Data Visual Instruction Data
VE & Input Size LLM Adapter ToP TuP Source Source
BLIP2 ViT-g/14 (EVA) & $224^{2}$ FlanT5-XL Q-Former+FC 3B 107M CC*-VG-SBU-L400 --
In-BL. ViT-g/14 (EVA) & $224^{2}$ Vicuna Q-Former+FC 7B 107M CC*-VG-SBU-L400 QA*
LA-V2 ViT-L/14 (CLIP) & $224^{2}$ LLaMA B-Tuning 7B 63.1M COCO Single-turn
LLaVA ViT-L/14 (CLIP) & $224^{2}$ Vicuna FC 7B 7B CC3M LLaVA-I
MGPT BLIP2-VE (EVA) & $224^{2}$ Vicuna FC 7B 3.1M CC-SBU-L400 CC+ChatGPT
m-owl ViT-L/14 (CLIP) & $224^{2}$ LLaMA LoRA+Q-Former 7B 388M CC*-CY-L400 LLaVA-I
Otter ViT-L/14 (CLIP) & $224^{2}$ LLaMA Resampler 9B 1.3B MIMIC-IT LLaVA-I
PGPT VIT-huge (ImageBind) & $224^{2}$ Vicuna Lora+FC 7B 28M -- LLaVA-I+CC+ChatGPT k
VPGT ViT-g/14 (CLIP) & $224^{2}$ Vicuna Q-Former 7B 107M COCO-VG-SBU-LC CC+ChatGPT
OF-2 ViT-L/14 (CLIP) & $224^{2}$ RedPajama Resampler 3B 63M LAION-2B, MMC4, ChatGPT ChatGPT
In-XC. ViT-g/14 (EVA)& $224^{2}$ internlm-xcomposer-7b Perceive Sampler+LoRA 7B 7B InternLM-XComposer-IT InternLM-XComposer-VI
L-1.5 ViT-L/14-336 (CLIP) & $336^{2}$ Vicuna FC 7B 7B LCS-558K LLaVA1.5-I
SGPT ViT-L/14-336 (CLIP) & $336^{2}$ Vicuna-v1.5 FC 7B 7.5B ShareGPT4V-PT ShareGPT4V+Vicuna-v1.5
Moe ViT-L/14-336 (CLIP) & $336^{2}$ Qwen-1.8B FC layer 2.2B 2.2B LCS-558K LLaVA1.5-I+others in  [7]
III-B Models
@@
We collect a total of 16 LVLMs, including 14 open-source models and 2 closed-source models, to create a model hub for evaluation. The open-source models consist of BLIP2 [32], LLaVA [51], MiniGPT-4 [33], mPLUG-owl [52], LLaMA-Adapter V2 [10], VPGTrans [53], Otter [4], InstructBLIP [28], PandaGPT [29], OpenFlamingo-V2 [31], InternLM-XComposer [5], LLaVA-1.5 [30], ShareGPT4V [6], and Moe-LLaVA [7]. The closed-source models are GeminiProVision [2] and GPT-4V(ision) [3]. To ensure a fair comparison, we carefully select the open-source LVLM versions with closely aligned parameter level. Model configurations and data configurations of the LVLMs are shown in Table I.
III-C Dataset
Our base dataset is derived from Tiny LVLM-eHub [27], which consists of 2,550 instructions and corresponding answers, and is organized into ten tasks that evaluate LVLM’s five multimodal capabilities:
Visual perception involves interpreting and understanding visual information, which is evaluated through tasks such as Image Classification, Object Counting (OC), and Multi-Class Identification (MCI) and includes a total of 450 images. Compared to Tiny LVLM-eHub [27], we add 50 additional images from the CIFAR-100 dataset [54].
Visual knowledge acquisition refers to the capability to acquire and understand visual information from images. This involves tasks such as Optical Character Recognition (OCR), Key Information Extraction (KIE), and image captioning and includes a total of 950 images. Compared to Tiny LVLM-eHub [27], we add 150 additional images from the following datasets (50 from each): POIE [55], MSCOCO_Caption_Karpathy [56], and WHOOPSCaption [57].
Visual reasoning involves the ability to reason and answer questions, which contains Visual Question Answering (VQA) and Knowledge-Grounded Image Description (KGID). There are a total of 750 images for this ability. Compared with Tiny LVLM-eHub [27], we include an additional 50 images from AOKVQAClose [58], 50 images from AOKVQAOpen [58], 50 images from WHOOPSWeird [57], and 50 images from Visdial [59].
Visual commonsense measures the model’s comprehension of shared human knowledge about visual concepts, which utilizes the same dataset as Tiny LVLM-eHub [27] and includes images related to color, shape, material, component, and other factors. A total of 250 images are used for assessing this ability.
Object hallucination refers to the phenomenon where LVLMs generate content that does not match the actual objects present in a given image, which uses the same dataset as Tiny LVLM-eHub [27] and includes 150 images.